# J1: Incentivizing Thinking in LLM-as-a-Judge via Reinforcement Learning

### v1

@@ Abstract
The progress of AI is bottlenecked by the quality of evaluation, and powerful LLM-as-a-Judge models have proved to be a core solution. Improved judgment ability is enabled by stronger chain-of-thought reasoning, motivating the need to find the best recipes for training such models to think. In this work we introduce J1, a reinforcement learning approach to training such models. Our method converts both verifiable and non-verifiable prompts to judgment tasks with verifiable rewards that incentivize thinking and mitigate judgment bias. In particular, our approach outperforms all other existing 8B or 70B models when trained at those sizes, including models distilled from DeepSeek-R1. J1 also outperforms o1-mini, and even R1 on some benchmarks, despite training a smaller model. We provide analysis and ablations comparing Pairwise-J1 vs Pointwise-J1 models, offline vs online training recipes, reward strategies, seed prompts, and variations in thought length and content. We find that our models make better judgments by learning to outline evaluation criteria, comparing against self-generated reference answers, and re-evaluating the correctness of model responses.
[…]
@@ 1 Introduction
Better judgments can be made by learning how to reason, which is observed in both humans and machines. For models, the ability to judge predictions is a vital process that is applied at all stages of development: during training and inference to provide a reward or verification signal, and during final benchmark evaluation to judge performance. Classical evaluation using reward models typically outputs a score directly (Ouyang et al., 2022) without having an explicit reasoning step. Using pre-trained and aligned language models to act as judges instead, i.e., LLM-as-a-Judge, allowed the possibility to generate chain-of-thought reasoning before making a decision, which was at first invoked by prompting (Zheng et al., 2023; Gu et al., 2024; Saha et al., 2024). Subsequently, iterative finetuning and direct preference optimization (DPO) methods were developed to improve these reasoning steps (Mahan et al., 2024; Wang et al., 2024d; Saha et al., 2025). In this work, we investigate recipes for further improvements to judgment reasoning via online Reinforcement Learning (RL).
Figure 1: Comparison of Pairwise-J1 with state-of-the-art LLM-as-a-Judge models on five reward modeling benchmarks at 8B and 70B scale. Pairwise-J1 outperforms all baselines despite being trained only on synthetic preference pairs.
@@
Our method, J1: Thinking-LLM-as-a-Judge via Reinforcement Learning, introduces a recipe for training such models. First, we convert the judgment task into a verifiable task for both verifiable (e.g., math problems) and typically subjective, non-verifiable prompts (e.g., user prompts from WildChat (Zhao et al., 2024)). This enables us to train a generalist judge across many types of tasks. To achieve this, we construct synthetic data for both categories of prompts by generating a high quality and a low quality response, such that pairwise judgment predictions are verifiable during training. We then train both the reasoning steps and judgment using Group Relative Policy Optimization Algorithm (GRPO;  Shao et al. 2024), utilizing a seed prompt and reward schemes designed to encourage thinking during judgment, in analogy to the approach in DeepSeek-R1 (Guo et al., 2025). Importantly, to mitigate positional bias, we introduce consistency-based rewards for pairwise judges and also train a pointwise judge (that is consistent by design) from pairwise supervision only. Both the reasoning process and final verdict or scores are then optimized during training.
@@
After running our best recipe at two model scales, J1-Llama-8B and J1-Llama-70B, we achieve strong performance across a variety of benchmarks: PPE (Frick et al., 2025), RewardBench (Lambert et al., 2024), JudgeBench (Tan et al., 2025), RM-Bench (Liu et al., 2025a), and FollowBenchEval (Saha et al., 2025). Our method outperforms: (i) models trained with SFT or Self-Taught reasoning (Wang et al., 2024d) and offline DPO (Saha et al., 2025); (ii) recent state-of-the-art generative reward models such as DeepSeek-GRM (Liu et al., 2025b) that is trained on significantly more data; (iii) models distilled from large reasoning models such as DeepSeek-R1; and (iv) closed reasoning models such as o1-mini (Jaech et al., 2024). While the much larger DeepSeek-R1 model still performs better on verifiable tasks, J1 achieves superior results on non-verifiable tasks. Additionally, test-time scaling of J1 leads to further improvements.
@@
We provide detailed ablations and analysis of these results by comparing our best recipe to other variants, that either modify the LLM-as-a-Judge setup (e.g., pairwise vs pointwise, with vs without scores), bias mitigation strategies, reward modeling approaches, or seed thinking prompt. In addition, we also analyze the distribution of the generated scores, thought lengths, and reasoning patterns within the thought generations. We observe that the model thinks about evaluation by outlining evaluation criteria, generating and comparing against reference answers, and re-evaluating the correctness of model responses (see Figure 2 for examples).
@@
Figure 2: Illustration of the thinking patterns of Pairwise-J1 and Pointwise-J1 models during RL training: it learns to outline evaluation criteria, generate reference answers, re-evaluate for correctness, and compare between responses. While Pairwise-J1 outputs a final verdict, indicating the better response, Pointwise-J1 generates a real-valued score, where a higher score indicates a better response.
@@ 2 J1: Thinking-LLM-as-a-Judge via Reinforcement Learning
2.1 Primary Problem Setup
Our primary evaluation setting is pairwise, in which an LLM judge takes an instruction $x$ and a pair of model responses $a,b$ as input and generates a preference judgment $y$, indicating the better response. The judge is conditioned on a seed thinking prompt (see Figure 7 in Appendix A) and generates intermediate thought tokens $t$, before deriving the final verdict. In the following subsections, we describe our training data and the Pairwise-J1 training recipe. Then in Section 2.5, we present alternative LLM-as-a-Judge settings and discuss how we derive RL recipes for them. An overview of our modeling setup is provided in Figure 3.
@@ 2.2 Synthetic Training Data Generation
The goal of J1 is to train a generalist judge that can evaluate LLM responses for both verifiable and non-verifiable tasks. With that motivation, we adopt the same synthetic training data used by Saha et al. (2025) to train EvalPlanner. Unlike most previous works (Liu et al., 2025b; Chen et al., 2025b), the use of synthetic data removes the reliance on human-annotated preference pairs since it can be costly to obtain (Wang et al., 2024a). It also ensures that J1 has a fair comparison with a prior state-of-the-art Thinking-LLM-as-a-Judge recipe which used DPO, allowing us to directly assess the benefits of our online RL-based training strategy.
J1’s final 22K training data consists of 17K WildChat (Zhao et al., 2024) and 5K MATH (Hendrycks et al., 2021) prompts, along with their corresponding preference pairs. For WildChat, rejected responses are obtained by prompting an LLM to first generate a “noisy” variant of the original instruction and then produce a response to this noisy instruction (Wang et al., 2024d). For MATH, rejected responses are sampled from generations by an LLM that do not lead to the gold answer. Given these preference pairs, we are thus able to convert the evaluation task into a verifiable task (i.e., predicting the better response), enabling the use of RL with verifiable rewards.
Pairwise LLM judges frequently suffer from a phenomenon called position bias, wherein the verdict changes if the order of the responses is swapped (Wang et al., 2024c; Chen et al., 2024). To mitigate this issue, we construct training data with both orders of response pairs – $(x,a,b)$ and $(x,b,a)$ – which are then fed into the corresponding thinking seed prompt (Figure 7 in Appendix A). Crucially, we construct position-agnostic batches of training data i.e., for the same input, we process both orders of responses in one batch (see Figure 3). In Section 2.3, we describe how this helps us design verdict consistency rewards for training J1 that can mitigate position bias in pairwise judges.
Figure 3: Reinforcement Learning recipes for training Pairwise-J1 and Pointwise-J1 models. We generate synthetic preference pairs for both verifiable and non-verifiable tasks to create position-agnostic training batches. Rewards based on verdict correctness, consistency, and score alignment jointly optimize evaluation thoughts and verdicts using online RL (GRPO). Pointwise-J1 is trained only via distant supervision from pairwise labels.
@@ 2.3 Reward Modeling
Following the training data construction, we describe J1’s reward modeling for RL training. We adopt a rule-based reward system that consists of two main types of rewards.
@@
Verdict Correctness Reward. J1’s primary reward is based on the accuracy of the final verdict. If a judge correctly predicts the better response, it receives a reward of $1$ and 00, otherwise.
Verdict Consistency Reward. To further encourage consistent reasoning and judgment, we introduce verdict (positional) consistency reward. In particular, we only assign a reward of $1$ when the model produces correct judgments for both input orders of the same response pair (i.e., $(x,a,b)$ and $(x,b,a)$).
Similar to DeepSeek-R1, we also experiment with format rewards to ensure that the thought tokens are enclosed in “<think>” and “</think>” tags, but do not observe noticeable improvements. In Section 4.2, we provide a detailed analysis of the effect of these reward types.
2.4 Primary J1 Recipe and Models
J1 jointly optimizes thoughts and judgments using GRPO, which makes RL training more efficient by removing the dependency on a separate critic model and instead estimating the baseline from group scores.
[…]
2.5 Other RL Training Recipes for Thinking-LLM-as-a-Judge
[…]
Models #Training PPE PPE PPE Correctness
Pref. Pairs Overall Preference Overall MMLU-Pro MATH GPQA MBPP-Plus IFEval
Open and Closed LLM-as-a-Judge
Llama-3.1-8B-Instruct 55.5 56.4 54.7 56.3 62.9 51.4 50.1 52.8
GPT-4o 62.3 67.1 57.6
Llama-3.3-70B-Instruct 65.8 65.9 65.7 72.1 73.1 61.2 59.6 62.3
SOTA Scalar Reward Models
Armo-8B-v0.1 1000K 60.9 60.6 61.2 66.0 71.0 57.0 54.0 58.0
Skywork-Reward-Gemma-2-27B 80K 55.6 56.6 54.7 55.0 46.2 44.7 69.1 58.3
DeepSeek-BTRM-27B 237K 67.5 68.3 66.7 68.8 73.2 56.8 68.8 66.0
SOTA Generative Reward Models
DeepSeek-GRM-27B 237K 62.2 64.7 59.8 64.8 68.8 55.6 50.1 59.8
DeepSeek-GRM-27B (MetaRM voting@32) 237K 65.2 67.2 63.2 68.1 70.0 56.9 50.8 70.4
EvalPlanner-Llama-8B 22K 54.1 55.5 52.8 57.0 59.0 50.3 47.7 50.0
EvalPlanner-Llama-70B 22K 67.9 65.6 70.2 78.4 81.7 64.4 62.2 64.3
J1-Llama-8B 22K 59.8 60.3 59.2 65.6 70.0 53.2 53.1 54.0
J1-Llama-8B (SC@32) 22K 61.3 60.6 61.9 67.5 76.6 55.7 54.6 54.9
J1-Llama-70B 22K 69.6 66.3 72.9 79.0 86.0 65.9 66.0 67.3
J1-Llama-70B (SC@32) 22K 70.4 67.0 73.7 79.9 88.1 66.5 66.5 67.2
Table 1: Results on PPE comparing Pairwise J1-Llama-8B (trained from Llama-3.1-8B-Instruct) and J1-Llama-70B (trained from Llama-3.3-70B-Instruct) with state-of-the-art LLM-as-a-Judge and reward models. $\dagger$: Results taken from Liu et al. (2025b) and Frick et al. (2025). $\ddagger$: Training data only contains synthetically constructed preference pairs.
@@ 3 Experimental Setup
Training. We implement J1 on top of verl (Sheng et al., 2024). All variants are trained on the 22K synthetic preference pairs described in Section 2.2. Appendix B includes other details of our experimental setup.
@@
  • RewardBench (Lambert et al., 2024) (3K samples), similar to JudgeBench, consists of preference pairs from 4 categories of prompts: chat, chat-hard, safety, and reasoning.
Consistent with prior work, we report accuracy over a single random ordering of paired responses for PPE, RewardBench, and RM-Bench. For JudgeBench and FollowBenchEval, we instead report position-consistent accuracy, where a sample is deemed correct only if the judge produces the correct verdict under both response orders. A more detailed analysis of positional consistency is presented in Section 4.2. Model selection is based on overall accuracy on RewardBench. Unless otherwise specified, inference is performed using vLLM (Kwon et al., 2023) with greedy decoding.
Baselines. We compare J1 to different categories of baselines: (i) LLMs acting as judges in a zero-shot manner (e.g., Llama-3.3-70B-Instruct, GPT-4o (Hurst et al., 2024)), (ii) state-of-the-art scalar reward models (e.g., DeepSeek-BTRM-27B (Liu et al., 2025b), Armo (Wang et al., 2024a), Skywork-Reward-Gemma-2-27B (Shiwen et al., 2024)), (iii) state-of-the-art generative reward models that belong to the same category as J1 (e.g., EvalPlanner (Saha et al., 2025), DeepSeek-GRM-27B (Liu et al., 2025b)), and (iv) general Reasoning/Thinking-LLMs (e.g., DeepSeek-R1-Distilled-Llama (Guo et al., 2025), DeepSeek-R1, and OpenAI-o1-mini).
Models PPE Preference PPE Correctness RewardBench RM-Bench JudgeBench FollowBenchEval
Llama-3.1-8B-Instruct 56.4 54.7 69.5 54.0 32.3 30.2
DeepSeek-R1-Distilled-Llama-8B 59.4 58.4 73.7 69.3 30.5 40.9
EvalPlanner-Llama-8B 55.5 53.0 83.0 68.1 30.2 45.3
J1-Llama-8B 60.3 59.2 85.7 73.4 42.0 48.3
Llama-3.3-70B-Instruct 65.9 65.7 85.4 69.5 48.6 52.2
DeepSeek-R1-Distilled-Llama-70B 66.1 71.1 86.9 80.8 46.0 54.6
EvalPlanner-Llama-70B 65.6 70.2 93.8 82.1 56.6 65.4
J1-Llama-70B 66.3 72.9 93.3 82.7 60.0 69.3
OpenAI-o1-mini 65.7 71.3 87.1 80.8 64.2 62.9
DeepSeek-R1-671B 68.0 76.5 90.6 88.6 68.9 71.7
Table 2: Results on five reward modeling benchmarks comparing Pairwise-J1 at 8B and 70B scale to a state-of-the-art Thinking-LLM-as-a-Judge model (EvalPlanner) and general Thinking-LLMs (o1-mini, distilled-R1, and R1). We report the default metric for each benchmark, where $\dagger$: JudgeBench and FollowBenchEval use positional consistent accuracy.

### v2

@@ Abstract
The progress of AI is bottlenecked by the quality of evaluation, making powerful LLM-as-a-Judge models a core solution. The efficacy of these judges depends on their chain-of-thought reasoning, creating a critical need for methods that can effectively optimize this reasoning process. In this work, we introduce J1, a reinforcement learning framework for teaching LLM judges to think before making decisions. Our core contribution lies in converting all judgment tasks for non-verifiable and verifiable prompts into a unified format with verifiable rewards, enabling direct optimization of evaluation quality while mitigating positional bias. We then use RL to train thinking-judges at scales of 8B, 32B, and 70B and show that they obtain state-of-the-art performance across multiple benchmarks. In particular, J1-Qwen-32B, our multitasked pointwise and pairwise judge also outperforms o1-mini, o3, and a much larger 671B DeepSeek-R1 on some benchmarks, while only training on synthetic data. Through comprehensive ablations of pairwise, pointwise, and multitask J1 variants, we demonstrate the effectiveness of our approach across seed prompts, reward strategies, and training recipes. Qualitative analysis reveals that J1 develops systematic evaluation strategies, including dynamic criteria generation, reference answer creation, iterative self-correction of initial assessments, and feedback generation for low-quality responses.
@@ 1 Introduction
Better judgments can be made by learning how to reason, which is observed in both humans and machines. For models, the ability to judge predictions is a vital process that is applied at all stages of development: during training and inference to provide a reward or verification signal, and during final benchmark evaluation to judge performance. Classical evaluation using reward models typically outputs a score directly (Ouyang et al., 2022) without having an explicit reasoning step. Using pre-trained and aligned language models to act as judges instead, i.e., LLM-as-a-Judge, allows the possibility to generate chain-of-thought reasoning before making a decision, which was at first invoked by prompting (Zheng et al., 2023; Gu et al., 2024; Saha et al., 2024). Subsequently, iterative finetuning and direct preference optimization (DPO) methods were developed to improve these reasoning steps (Mahan et al., 2024; Wang et al., 2024d; Yu et al., 2025a; Saha et al., 2025). In this work, we investigate approaches for further improvements to judgment reasoning via online Reinforcement Learning (RL).
@@
We introduce J1 (Thinking-LLM-as-a-Judge via RL), a framework that incentivizes LLMs to think for evaluation through three primary aspects: (i) Unified Verifiable Training: we design a unified training recipe that converts all judgment tasks, from both verifiable (e.g., math problems) and typically subjective, non-verifiable prompts (e.g., user prompts from WildChat (Zhao et al., 2024)), into a format that can be optimized with RL from verifiable rewards. This allows us to train a single, generalist judge across diverse domains using only synthetic data. (ii) Reasoning-Optimized Training: we use GRPO (Shao et al., 2024) to directly optimize the quality of evaluation thoughts, in analogy to the approach in DeepSeek-R1 (Guo et al., 2025a). Guided by a seed prompt and targeted reward schemes, J1 teaches the model to reason critically about evaluation. (iii) Multitask and Bias-Aware Judge: we address positional bias through consistency-based rewards and, more importantly, develop a method to train inherently consistent pointwise judges using only pairwise supervision. We then unify these approaches into a single multitask model that is capable of performing both pointwise and pairwise evaluations.
@@
After applying our best J1 training recipe, we train judges on top of Llama-3.1-Instruct (Grattafiori et al., 2024) and Qwen3 (Yang et al., 2025) models at 8B, 32B, and 70B scales that achieve state-of-the-art performance across a variety of benchmarks: PPE (Frick et al., 2025), RewardBench (Lambert et al., 2025), JudgeBench (Tan et al., 2025), RM-Bench (Liu et al., 2025a), and FollowBenchEval (Saha et al., 2025). In particular, J1 significantly outperforms: (i) models trained with SFT or Self-Taught reasoning (Wang et al., 2024d) and offline DPO (Saha et al., 2025); (ii) scalar RMs trained with a Bradley-Terry objective such as Skywork-RM (Shiwen et al., 2024); (iii) recent state-of-the-art generative reward models such as DeepSeek-GRM (Liu et al., 2025b) and Reasoning Reward Model (Guo et al., 2025b) that are trained on significantly more data; (iv) open-weight reasoning models like DeepSeek-R1-671B; and (v) closed reasoning models such as o1-mini (Jaech et al., 2024) and o3 (OpenAI, 2025). Crucially, J1 achieves these results while only training a 32B model and leveraging a small amount of synthetic training data. Test-time scaling of J1 via majority vote over multiple verdicts or averaging over multiple judgment scores leads to further improvements.
@@
We provide detailed ablations and analysis of these results by comparing our best training strategy to other variants, that either modify the LLM-as-a-Judge setup (e.g., pairwise vs pointwise vs joint/multitasked, with vs without scores), bias mitigation strategies, reward modeling approaches, or seed thinking prompt. We show that our proposed joint model flexibly performs both pointwise and pairwise evaluations, while also being superior to both separately trained counterparts. In addition, we also analyze the distribution of the generated scores, thought lengths, and reasoning patterns within the thought generations. Our qualitative analysis shows that J1 models learn to make better judgments by systematically outlining evaluation criteria, comparing responses against self-generated reference answers, critically re-evaluating their own initial assessments, and providing feedback (see Figures 167, and 8 for examples).
@@
Figure 1: Illustration of the thinking patterns of pairwise and pointwise J1 models during RL training. J1 learns to outline evaluation criteria, generate reference answers, re-evaluate correctness, and compare between responses. Pairwise setup outputs a final verdict indicating the better response, pointwise generates a real-valued score, with a higher score for the better response. More examples of J1’s thinking traces are shown in Figures 67, and 8.
@@ 2 J1: Thinking-LLM-as-a-Judge via Reinforcement Learning
Our J1 framework trains an LLM-as-a-Judge that generates chain-of-thought reasoning before scoring responses or making a preference judgment. Our primary setting is pairwise evaluation, where the judge takes an instruction $x$ and two responses $(a,b)$ to produce a verdict $y$ indicating the preferred response. The model’s output consists of intermediate thought tokens $t$ followed by the final verdict $y$, conditioned on a seed prompt designed to elicit reasoning (see Appendix B).
This section details the three core components of our method. First, we describe the synthetic data generation process that creates verifiable tasks for RL training (§2.1). Next, we define the reward functions used to optimize for both judgment correctness and consistency (§2.2). Finally, we outline different judge formulations and training, including pairwise, pointwise, and multitask model (§2.3). An overview of the overall process is shown in Figure 2.
@@ 2.1 Synthetic Training Data Generation
The goal of J1 is to train a generalist judge for both verifiable and non-verifiable tasks. To achieve this, we follow the unified training dataset of synthetic preference pairs from Wang et al. (2024a), which removes the reliance on costly human annotations used in prior work (Liu et al., 2025b; Chen et al., 2025b). By building on the same data generation strategy from Saha et al. (2025), we can directly compare the effectiveness of our online RL approach against prior work using offline DPO on the same data distribution.
Our final 22K training data consists of 17K WildChat (Zhao et al., 2024) and 5K MATH (Hendrycks et al., 2021) prompts, along with their corresponding preference pairs. For WildChat, rejected responses are obtained by prompting an LLM to first generate a “noisy” variant of the original instruction and then produce a response to this noisy instruction (Wang et al., 2024d). See Figure 14 for the prompt and Table 5 for an example of a synthetically generated training pair. For MATH, rejected responses are sampled from generations by an LLM that do not lead to the gold answer. Given these preference pairs, we are thus able to convert the evaluation task into a verifiable task (i.e., predicting the better response), enabling the use of RL with verifiable rewards.
To address the known issue of position bias in pairwise judges (Wang et al., 2024c; Chen et al., 2024), we augment the training data by including both response orderings – $(x,a,b)$ and $(x,b,a)$, using the thinking seed prompt (Figure 9 in Appendix B). We construct training batches to be position-agnostic, i.e., both orderings of a given pair are processed in the same batch. As detailed in the next section, this batching strategy is crucial for implementing our consistency-based rewards.
@@ 2.2 Reward Modeling
We adopt a straightforward and effective rule-based reward system designed to encourage accurate and consistent judgments.
@@
Verdict Correctness. J1’s primary reward signal is binary. The model receives a reward of +1+1 if its final verdict correctly identifies the preferred response, and 0 otherwise.
Verdict Consistency. To explicitly mitigate position bias, we introduce a consistency reward. A reward of +1+1 is granted only if the model produces the correct verdict for both input orderings of a response pair, i.e., for $(x,a,b)$ and $(x,b,a)$. An incorrect verdict on either ordering results in a reward of 0.
We also explored adding format-based rewards to enforce the use of “<think>” tags, but found no noticeable performance benefit. The effects of different reward components are ablated in §4.2.
Figure 2: Reinforcement Learning recipes for training J1 models. We generate synthetic preference pairs for both verifiable and non-verifiable tasks to create position-agnostic training batches. Rewards based on verdict correctness, consistency, and score alignment jointly optimize evaluation thoughts and verdicts using online GRPO. Pointwise-J1 is trained only via distant supervision from pairwise labels. MultiTask-J1 combines pairwise and pointwise formulation.
2.3 J1 Formulations and Training
Having described the training data and reward schemes, we now use GRPO (Shao et al., 2024) to jointly optimize thought generation and final judgments. We explore several training formulations that differ in their input/output formats and reward schemes.
Pairwise J1 with Verdict (PaV). Our first formulation, $\texttt{J1}{}_{\texttt{PaV}}:\texttt{prompt}_{\texttt{PaV}}(x,a,b)\rightarrow(t,y)$, receives a user question and a response pair, and generates thought tokens and the preferred response (as the final verdict). Figure 9 in Appendix B shows the seed thinking prompt. Figure 1 illustrates examples of judgment generation with this formulation.
Pairwise J1 with Scores (PaS). Instead of directly generating a verdict, our pairwise score-based variant $\texttt{J1}{}_{\texttt{PaS}}:\texttt{prompt}_{\texttt{PaS}}(x,a,b)\rightarrow(t,s_{a},s_{b})$, generates real-valued scores $s_{a}$, $s_{b}$ for response $a$ and $b$, respectively. The response that obtains the higher score is chosen as the final verdict. Figure 11 in Appendix B shows the corresponding thinking prompt. To train a model with this recipe, we replace the verdict-based reward with a score-based reward: a binary reward is assigned depending on whether the predicted scores are consistent with the gold verdict.
Pairwise J1 with Scores&Verdict (PaVS). In another pairwise variant $\texttt{J1}{}_{\texttt{PaVS}}:\texttt{prompt}_{\texttt{PaVS}}(x,a,b)\rightarrow(t,s_{a},s_{b},y)$ , the model generates scores for both responses as well as the final verdict, where the generated scores are interpreted as observations of unknown latent variables to help with the pairwise judgment task; therefore $y$ is directly used as the final verdict. Consequently, the reward is also computed using only the final verdict (and not the intermediate scores). Figure 12 in Appendix B shows the corresponding thinking prompt.
Pointwise J1 (PoS). We also introduce a pointwise judge, $\texttt{J1}{}_{\texttt{PoS}}:\texttt{prompt}_{\texttt{PoS}}(x,a)\rightarrow(t,s)$, which takes an instruction $x$ and a single response $a$ as input, and outputs a score $s$ that reflects the quality or reward of the response. Unlike pairwise judges, pointwise judges are inherently consistent. We train $\texttt{J1}{}_{\texttt{PoS}}$ via distant supervision with the same pairwise training data as used for all our pairwise variants. Each preference pair is split into two separate pointwise samples and the model is trained with a seed thinking prompt that assigns a score between 0 and 10 to a given response (see Figure 13 in Appendix B). Both scores are evaluated jointly, and the model receives a reward of 1 if the scores are consistent with the gold verdict. Since preference rankings are significantly easier to obtain than pointwise annotations, Pointwise-Thinking-LLM-Judge represents one of our novel contributions.
Multitask Pairwise & Pointwise J1 (MT). Finally, we unify the Pointwise and Pairwise (PaS) paradigms into a single multitask model by jointly training on both pairwise and pointwise data. This model has the advantage of being a unified model that can be evaluated in both setups and improves upon separately trained judges. As we show later, given the overall superiority of pairwise judgments over pointwise judgments, we evaluate the multitasked model in a pairwise setup for best results.
@@ 3 Experimental Setup
Training. We implement J1 on top of verl (Sheng et al., 2024). All variants are trained on the 22K synthetic preference pairs described in Section˜2.1. Appendix C includes other details of our experimental setup.
@@
  • RewardBench (Lambert et al., 2025) (3K samples), similar to JudgeBench, consists of preference pairs from 4 categories of prompts: chat, chat-hard, safety, and reasoning.
Consistent with prior work, we report accuracy over a random ordering of paired responses for PPE, RewardBench, and RM-Bench. For JudgeBench and FollowBenchEval, we instead report position-consistent accuracy, where a sample is deemed correct only if the judge produces the correct verdict under both response orders. A more detailed analysis of position consistency is presented in Section˜4.2. Model selection is based on overall accuracy on RewardBench. Inference is performed using vLLM (Kwon et al., 2023).
Baselines. We compare J1 to different categories of baselines: (i) non-thinking LLMs acting as judges in a zero-shot manner (e.g., Llama-3.3-70B-Instruct, GPT-4o (Hurst et al., 2024), (ii) thinking LLMs acting as judges (e.g., DeepSeek-R1-Distilled-Llama (Guo et al., 2025a), DeepSeek-R1, Qwen3-32B (Yang et al., 2025), OpenAI-o1-mini, o3 (OpenAI, 2025)) (iii) state-of-the-art scalar reward models (e.g., DeepSeek-BTRM-27B (Liu et al., 2025b), Armo (Wang et al., 2024a), Skywork-Reward-Gemma-2-27B (Shiwen et al., 2024)), (iv) state-of-the-art generative reward models that belong to the same category as J1 (e.g., EvalPlanner (Saha et al., 2025), DeepSeek-GRM-27B (Liu et al., 2025b)), and Reasoning Reward Model (Guo et al., 2025b). Additionally, for RewardBench, we compare J1 to all highest-ranked Generative Reward Models according to the leaderboard. [^1] https://huggingface.co/spaces/allenai/reward-bench
Table 1: Results on PPE Correctness comparing J1 against other LLM-Judges and reward models. All models are evaluated in a pairwise setup. $\dagger$: Results from Liu et al. (2025b) and Frick et al. (2025). $\ddagger$: Trained solely on synthetically constructed preference pairs. Reported gains are relative to the corresponding base models.
PPE Correctness #Training Pref. Pairs Overall MMLU-Pro MATH GPQA MBPP-Plus IFEval
Base LLM-as-a-Judge
Llama-3.1-8B-Instruct 54.7 00056.3 62.9 51.4 50.1 52.8
Llama-3.3-70B-Instruct 65.7 00072.1 73.1 61.2 59.6 62.3
Qwen3-32B (thinking) 66.5 00075.6 85.2 53.6 55.9 62.0
SOTA Scalar Reward Models
Armo-8B-v0.1 1000K 61.2 00066.0 71.0 57.0 54.0 58.0
Skywork-Reward-Gemma-2-27B 0080K 54.7 00055.0 46.2 44.7 69.1 58.3
DeepSeek-BTRM-27B 0237K 66.7 00068.8 73.2 56.8 68.8 66.0
SOTA Generative Reward Models
DeepSeek-GRM-27B 0237K 59.8 00064.8 68.8 55.6 50.1 59.8
EvalPlanner-Llama-8B 0022K 52.8 00057.0 59.0 50.3 47.7 50.0
EvalPlanner-Llama-70B 0022K 70.2 00078.4 81.7 64.4 62.2 64.3
J1 Models (Ours)
J1-Llama-8B 0022K 59.2 +4.5 00065.6 +9.3 70.0 +7.1 53.2 +1.8 53.1 +3.0 54.0 +1.2
J1-Llama-70B 0022K 72.9 +7.2 00079.0 +6.9 86.0 +12.9 65.9 +4.7 66.0 +6.4 67.3 +5.0
J1-Qwen-32B 0022K 74.6 +8.1 00082.2 +6.6 93.3 +8.1 65.2 +11.6 65.3 +9.4 66.8 +4.8
J1-Qwen-32B-MultiTask 0022K 76.8+10.3 00085.0+9.4 94.3+9.1 68.6+15.0 66.3 +10.4 69.5+7.5